# Best Arm Identification for Prompt Learning under a Limited Budget

### v2

@@ 3.2 A Bandit View of Prompt Learning
  • Using a prompt $p$ to interact with an LLM can be viewed as selecting a bandit arm $k$ to pull in the MAB model;
@@
It can be further recognized that the target of prompt learning is more suitable to be captured as the best arm identification (BAI) problem, instead of a regret minimization one, as it only cares about finding the optimal prompt $p^{*}$ instead of the interaction performed during the learning process.
With the relationship between prompt learning and BAI established, we further consider the constraint of learning under a limited budget. We argue that this aligns with one of the main research directions in BAI called fixed-budget best arm identification (BAI-FB) (Karnin et al.,, 2013; Wang et al.,, 2023; Gabillon et al.,, 2012). BAI-FB particularly considers the problem of
maximizing the probability of correctly identifying the best arm $k^{*}$ while not pulling arms more than $T$ times.
It can be observed that this formulation matches the goal of prompt learning under a limited budget; thus BAI-FB provides us with a perfect toolbox to enhance the commonly required prompt selection process. The connection between prompt learning and MAB, in particular, BAI-FB, is further illustrated in Table 1. To avoid confusion, in the remainder of this paper, we will adopt the notation of prompt learning as introduced in Sec. 2.
Table 1: Connections between Prompt Learning and MAB.
Prompt Learning Multi-armed Bandits
@@ 3.3 Harnessing the Power of BAI-FB
As mentioned, we recognize that prompt learning under a limited budget is a matching application scenario for BAI-FB. In this paper, we propose a general framework called TRIPLE (besT aRm Identification for Prompt LEarning) to harness the power of BAI-FB in solving the prompt learning problem. This is possible because BAI-FB has witnessed significant development over the years, with several efficient designs being proposed. As a first step, we choose two popular and successful BAI-FB schemes and implement them for prompt learning, which are briefly described below.
Sequential Halving (SH) (Karnin et al.,, 2013): SH is one of the first provably efficient BAI-FB designs and remains popular after a decade of its proposal. It follows a protocol that divides the total budget $N$ into $\lceil\log_{2}(|{\mathcal{P}}|)\rceil$ equal-length phases. In each phase, SH uniformly tries all active prompts (initialized as ${\mathcal{P}}$) and eliminates half of them with the lower sample means for the next phase. The final active arm is output as the identified optimal prompt. A complete description of TRIPLE-SH is provided in Alg. 2 of Appendix A.
Continuously Reject (CR) (Wang et al.,, 2023): CR is a recently proposed method, which can be viewed as an extension of the classical Successively Reject (SR) design (Audibert et al.,, 2010). It uniformly explores active prompts (initialized as ${\mathcal{P}}$) and performs potential elimination of poorly-performed prompts after each pull. The elimination is based on carefully designed criteria using the Large Deviation Principle. It can be observed that, without the phased structure, CR is more adaptive than SH (and SR), which makes it appealing both theoretically and practically. We present details of TRIPLE-CR in Alg. 3 of Appendix A.
While MAB has found broad applications in recommender systems (Li et al.,, 2010), healthcare (Shen et al.,, 2020), wireless communications (Gai et al.,, 2012), and beyond (Bouneffouf and Rish,, 2019), a systematical connection between MAB and prompt learning has not been established before to the best of our knowledge, which may spark new research activities (see discussions in Sec. 7). In addition, although SH and CR are selected as the representatives, the connection between prompt learning and MAB is fundamental, e.g., any existing or forthcoming BAI-FB designs can be flexibly incorporated into TRIPLE.
@@ 4 Handling Large Candidate Pools via Prompt Embeddings
The connection established in the last section provides us with the core idea of leveraging BAI-FB designs to tackle prompt learning. As having been theoretically established (Audibert et al.,, 2010), solving BAI-FB without additional structures, however, will unavoidably incur an identification error that is positively related to the number of candidate prompts $|{\mathcal{P}}|$. In other words, given a larger pool of prompts, it becomes harder to find the optimal prompt with the basic BAI-FB designs, which restricts their applicability to practical prompt learning problems.
The key reason behind this is that each candidate prompt is treated independently in the basic BAI-FB. Thus, budgets need to be assigned to all of the prompts and no information can be shared among them, which is often not the case in prompt learning. For a prompt learning problem, the underlying task is often stable, e.g., rewriting emails, constructing TLDR, etc. The candidate prompts, regardless of their generation methods, should all reflect the purpose of the underlying task and thus share similarities. For example, the candidate prompts generated via demonstrating LLMs (Zhou et al.,, 2022) often share similar structures and differ only in a few words or word orders.
With the above observation, we target sharing information among prompts during learning. To achieve this, we propose to leverage an embedding model, denoted as $\texttt{embed}:{\mathcal{V}}\to{\mathbb{R}}^{d}$, to obtain the sentence embedding of the prompts:
$e(p):=\texttt{embed}(p)\in{\mathbb{R}}^{d},~{}~{}{\mathcal{E}}:=\{e(p):p\in{\mathcal{P}}\},$
where $d$ refers to the embedding dimension. In the experiments, the OpenAI embedding API is adopted for embedding while, in general, any sufficiently expressive models can be incorporated. Also, due to this flexibility, using embedding models is fundamentally different from requiring a white-box LLM as in Chen et al., (2023); Lin et al., (2023).
With the obtained prompt embeddings, we propose two useful enhancements to further improve the learning effectiveness when the pool of candidate prompts is large.
@@
4.1 Leveraging Similarities via Clustering
Since the key challenge is a large pool of candidate prompts, an intuitive idea is to effectively decrease the size of the pool. We thus propose a two-phased BAI-FB scheme for prompt learning. In Phase I, the entire pool of candidate prompts is clustered into several groups based on their embeddings, and BAI-FB is performed on the clusters with an initial target of finding the optimal cluster (or the few good clusters). Then, in Phase II, BAI-FB is performed on the prompts in the optimal cluster with the target of identifying one final prompt. For both phases, different BAI-FB designs can be incorporated, e.g., SH and CR. The entire procedure, referred to as TRIPLE-CLST, is described in Alg. 1.
The effectiveness of TRIPLE-CLST relies on the clustering results produced in Phase I. Ideally, prompts with similar performances should be clustered together. In this way, Phase I can quickly eliminate the prompts with poor performances, leaving a small pool of good prompts for Phase II to process. In the experiments, this intuitive phenomenon is indeed observed. Here we provide examples in Fig. LABEL:fig:cluster (with more details in Appendix C.1), where as expected, prompts in the same cluster share similar performances.
@@ 4.2 Sharing Information via Function Approximation
Besides clustering, another idea to incorporate the prompt features is to learn a common function (e.g., an MLP) to predict the prompt features based on their features. Similar ideas of function approximation have also been widely adopted in MAB literature to share information among large action spaces, with function classes ranging from linear functions (Abbasi-Yadkori et al.,, 2011; Yang and Tan,, 2022) to neural networks (Zhu et al.,, 2021; Zhou et al.,, 2020).
In the setting considered in this work, we adopt a recently developed BAI-FB scheme as described in the following:
GSE (Azizi et al.,, 2023): It inherits the general phased elimination flow of SH described in 3.3. The major difference is that SH uses sample means to perform eliminations. GSE, on the other hand, leverages collected samples from previous phases to train a reward function $g_{\bm{\theta}}(\cdot):{\mathbb{R}}^{d}\to{\mathbb{R}}$ that maps prompt embeddings to the predicted performance, which is further used to eliminate prompts. The details of TRIPLE-GSE are provided in Alg. 4 of Appendix A.
@@ 5 Experiments
In this section, extensive experimental results are reported to evaluate the efficiency of TRIPLE across diverse prompting tasks from two standard datasets: Instruction-Induction (Honovich et al.,, 2022) and BigBench (Suzgun et al.,, 2022). The results reported in this section are collected from gpt-3.5-turbo-1106 model of OpenAI (OpenAI, 2023a, ) as well as Llama2-7b (Touvron et al., 2023b, ). Full experimental details can be found in Appendix B. We particularly focus on $12$ representative tasks, while the complete results of the overall $47$ tasks are reported in Appendix C.
@@ 5.1 Evaluating TRIPLE with Fixed Prompt Pools
As TRIPLE main focuses on the prompt selection component, we perform initial evaluations in an isolated fashion of selecting from fixed pools of candidate prompts. For this experiment, candidate pools of prompts are generated following the well-established APE design (Zhou et al.,, 2022) with a high LLM temperature to ensure randomness. Then, under a limited budget, the performances of TRIPLE algorithms are compared with the following two baselines:
@@
  • Uniform. Many previous designs choose to evaluate the entire candidate pool on all development data, e.g., Guo et al., (2023); Prasad et al., (2022). With a limited budget, this strategy corresponds to a uniform allocation, i.e., equally dividing the total budget to test all prompts.
@@
  • UCB. The upper confidence bound (UCB) method is a famous design for regret minimization in MAB. Its variants have been applied to prompt selection in Pryzant et al., (2023); Lin et al., (2023). We evaluate UCB using its vanilla version from Auer et al., (2002).
Performance with small prompt pools. We begin with relatively small candidate pools, generating $10$ prompts for selection in each task. Results with an overall budget of $50$ are reported in Fig. LABEL:fig:bar_fig_main (top). It can be observed that TRIPLE methods outperform two baselines on most of the tasks, evidencing its prowess in allocating limited budgets for prompt selection. Especially, across these tasks, TRIPLE-CR consistently outperforms Uniform with improvements ranging from $8\%$ to $129\%$, and on average, a notable gain of $35\%$. Moreover, it achieves a $3\%$ to $30\%$ improvement over UCB, with an average $13\%$ gain. The benefit of using TRIPLE over Uniform is intuitive as BAI-FB designs are much more adaptive. The improvement over UCB further highlights the ineffectiveness of methods for regret minimization in prompt learning and provides empirical justifications for using BAI-FB.
@@
Performance with large prompt pools. We next test larger candidate pools with $30$ prompts per task. Results reported in Fig. LABEL:fig:bar_fig_main (middle) reflect the selection performance with an overall budget of $150$. It can be observed that while TRIPLE-SH and TRIPLE-CR still achieve better performance than Uniform and UCB, the enhanced TRIPLE-CLST and TRIPLE-GSE demonstrate more remarkable improvements, evidencing their capabilities in leveraging prompt embeddings to handle large prompt pools. In particular, both TRIPLE-CLST and TRIPLE-GSE significantly outperform the previously best TRIPLE-CR method (let alone Uniform and UCB), with a gain ranging from $2\%$ to $60\%$, averaging of $13\%$ and $16\%$, respectively. With the same prompt pools, we also test the performance on Llama2-7b with results shown in Fig. LABEL:fig:bar_fig_main (bottom). We can see that the superiority of TRIPLE is even more evident on Llama-2: TRIPLE-GSE and TRIPLE-CLST achieve average gains of $45\%$ and $51\%$ over Uniform across all tasks, reaching a remarkable $200\%$ improvement on the “larger__\__animal” task.
@@
Impact of the total budget. For a more comprehensive understanding, using candidate pools with $30$ prompts, we further examine the impact of budgets, starting with $5$ evaluations per prompt on average (i.e., $150$ overall as adopted in Fig. LABEL:fig:bar_fig_main (middle)), and then gradually increasing to $30$ (i.e., $900$ overall, which is the same as the experiments in Zhou et al., (2022)). From the results shown in Fig. LABEL:fig:budget (left), we see that the improvements of TRIPLE over baselines are more pronounced with lower budgets. In particular, with a budget of $10$ evaluations per prompt on average, TRIPLE-CR, TRIPLE-CLST and TRIPLE-GSE maintain notable $9.7\%$, $13.5\%$ and $17.4\%$ improvement over Uniform, respectively; when the budget escalates to $20$ evaluations per prompt on average, TRIPLE-CLST and TRIPLE-GSE still achieve an approximate $8\%$ improvement. Once the budget reaches $30$ evaluations per prompt on average, all methods provide approximately the same performance as they can all identify the optimal prompts under this generous budget.
To further guide practical implementation, we additionally investigate how to select a reasonable budget. In particular, we focus on the efficiency of various prompt selection algorithms in identifying a “good” prompt – either the optimal prompt in the pool or achieving $95\%$ or better of the optimal prompt’s performance. Fig. LABEL:fig:budget (right) illustrates that initial increases in budgets significantly improve the probability of identifying a good prompt, but this benefit tapers off with further budget expansion. This finding suggests that starting with a modest budget and incrementally increasing it is the more effective approach, stopping when additional investment no longer translates into significant returns.
@@
Table 2: Performances of integrating TRIPLE in the end-to-end pipelines. The baseline methods reported in the original implementations are labeled as (b). For each task, the best score across two pipelines is marked as red, and the best score in the remaining pipeline is highlighted as yellow. TRIPLE-CR are selected over TRIPLE-SH due to its better performance observed in the previous experiments. TRILE-CLST is ignored in the tests with APO, as it is ineffective to cluster only $10$ prompts.

### v3

@@ 3.2 A Bandit View of Prompt Optimization
  • Using a prompt $p$ to interact with LLM can be viewed as selecting a bandit arm $k$ to pull in MAB;
@@
It can be further recognized that the target of prompt optimization is more suitable to be captured as the best arm identification (BAI) problem, instead of a regret minimization one, as it only cares about finding the optimal prompt $p^{*}$ instead of the cumulative performance of interactions performed during the learning process.
Table 1: Prompt Optimization and MAB.
Prompt Optimization Multi-armed Bandits
@@
With the relationship between prompt optimization and BAI established, we further consider the constraint of learning under a limited budget. We argue that this aligns with one of the main research directions in BAI called fixed-budget best arm identification (BAI-FB) (Karnin et al.,, 2013; Wang et al.,, 2023; Gabillon et al.,, 2012). BAI-FB particularly considers the problem of maximizing the probability of correctly identifying the best arm $k^{*}$ while not pulling arms more than $T$ times. It can be observed that this formulation matches the goal of prompt optimization under a limited budget; thus BAI-FB provides a perfect toolbox to enhance the commonly required prompt selection process. The connection between prompt optimization and MAB, in particular, BAI-FB, is further illustrated in Table 1. To avoid confusion, in the remainder of this paper, we will adopt the notation of prompt optimization as introduced in Sec. 2.
@@ 3.3 Harnessing the Power of BAI-FB
As mentioned, we recognize that prompt optimization under a limited budget is a matching application scenario for BAI-FB. In this paper, we propose a general framework called TRIPLE (besT aRm Identification for Prompt LEarning) to harness the power of BAI-FB in solving the prompt optimization problem. This is possible because BAI-FB has witnessed significant development over the years, with several efficient designs being proposed. As a first step, we choose two popular and successful BAI-FB schemes and implement them for prompt optimization, which are briefly described below. Their complete descriptions are provided in Algs. 2 and 3 of Appendix C.
Sequential Halving (SH). SH is one of the first provably efficient BAI-FB designs (Karnin et al.,, 2013) and remains popular after a decade of its proposal. It follows a protocol that divides the total budget $N$ into $\lceil\log_{2}(|{\mathcal{P}}|)\rceil$ equal-length phases. In each phase, SH uniformly tries all active prompts (initialized as ${\mathcal{P}}$) and eliminates half of them with the lower sample means for the next phase. The final active arm is output as the identified optimal prompt.
Continuously Reject (CR). CR is a recently proposed method (Wang et al.,, 2023), which can be viewed as an extension of the classical Successively Reject (SR) design (Audibert et al.,, 2010). It uniformly explores active prompts (initialized as ${\mathcal{P}}$) and performs potential elimination of poorly-performed prompts after each pull. The elimination is based on carefully designed criteria using the Large Deviation Principle. It can be observed that, without the phased structure, CR is more adaptive than SH (and SR), which makes it appealing both theoretically and practically.
While MAB has found broad applications in recommender systems (Li et al.,, 2010), healthcare (Shen et al.,, 2020), wireless communications (Gai et al.,, 2012), and beyond (Bouneffouf and Rish,, 2019), a systematical connection between MAB and prompt optimization has not been established before to the best of our knowledge, which may spark new research activities (see discussions in Sec. 7). In addition, although SH and CR are selected as the representatives, the connection between prompt optimization and MAB is fundamental. Any existing or forthcoming BAI-FB designs can be flexibly incorporated into TRIPLE, e.g., the Bayesian perspective provided in Komiyama et al., (2023); Atsidakou et al., (2022)
Remark 3.1.
As mentioned in Sec. 1, Pryzant et al., (2023); Lin et al., (2023) leverage specific MAB designs to perform prompt selection without a comprehensive discussion as above on their connection. Moreover, Pryzant et al., (2023) argues that UCB (Auer et al.,, 2002) is suitable, while Lin et al., (2023) also uses a UCB-variant, NeuralUCB (Zhou et al.,, 2020), as the core method. However, both of UCB and NeuralUCB are designed for regret minimization (i.e., optimizing the cumulative interaction performance during learning). As illustrated in Sec. 3.1, designs for regret minimization cannot achieve optimal performance for the goal of identifying the optimal arm (i.e., BAI), which thus are not well-suited for prompt optimization.
@@ 4 Handling Large Candidate Pools via Prompt Embeddings
The connection bulit in the last section provides us with the core idea of leveraging BAI-FB designs to tackle prompt optimization. As having been theoretically established (Audibert et al.,, 2010), solving BAI-FB without additional structures, however, will unavoidably incur an identification error that is positively related to the number of candidate prompts $|{\mathcal{P}}|$. In other words, given a larger pool of prompts, it becomes harder to find the optimal prompt with the basic BAI-FB designs, which restricts their applicability to practical prompt optimization problems (where possibly the number of prompts exceeds the budget).
The key reason behind this is that each candidate prompt is treated independently in the basic BAI-FB. Thus, budgets need to be assigned to all the prompts and no information can be shared among them, which is often not the case in prompt optimization. For a prompt optimization problem, the underlying task is often stable, e.g., rewriting emails, constructing TLDR, etc. The candidate prompts, regardless of their generation methods, should all reflect the purpose of the underlying task and thus share similarities. For example, the candidate prompts generated via demonstrating LLMs (Zhou et al.,, 2022) often share similar structures and differ only in a few words or word orders.
With the above observation, we target sharing information among prompts during learning. To achieve this, we propose to leverage an embedding model, denoted as $\texttt{embed}:{\mathcal{V}}\to{\mathbb{R}}^{d}$, to obtain the sentence embedding of the prompts: $e(p):=\texttt{embed}(p)\in{\mathbb{R}}^{d},~{}~{}{\mathcal{E}}:=\{e(p):p\in{\mathcal{P}}\}$, where $d$ refers to the embedding dimension. In the experiments, the OpenAI embedding API is adopted while, in general, any sufficiently expressive models can be incorporated. Also, due to this flexibility, using embedding models is fundamentally different from requiring a white-box LLM (Chen et al.,, 2023; Lin et al.,, 2023). With the obtained prompt embeddings, we propose two useful enhancements to further improve the learning effectiveness when the pool of candidate prompts is large.
4.1 Leveraging Similarities via Clustering
@@
Since the key challenge is a large pool of candidate prompts, an intuitive idea is to effectively decrease the size of the pool. We thus propose a two-phased BAI-FB scheme for prompt optimization. In Phase I, the entire pool of candidate prompts is clustered into several groups based on their embeddings, and BAI-FB is performed on the clusters with an initial target of finding the optimal cluster (or the few good clusters). Then, in Phase II, BAI-FB is performed on the prompts in the optimal cluster with the target of identifying one final prompt. For both phases, different BAI-FB designs can be incorporated, e.g., SH and CR. The entire procedure, referred to as TRIPLE-CLST, is described in Alg. 1.
The effectiveness of TRIPLE-CLST relies on the clustering results produced in Phase I. Ideally, prompts with similar performances should be clustered together. Then, Phase I can quickly eliminate the prompts with poor performances, leaving a small pool of good prompts for Phase II to process. In the experiments, this intuitive phenomenon is indeed observed. In particular, in Fig. LABEL:fig:cluster of Appendix F.1, as expected, prompts in the same cluster share similar performances.
@@ 4.2 Sharing Information via Function Approximation
Besides clustering, another idea to incorporate the prompt embeddings is to learn a common function (e.g., an MLP) to predict the prompt performances based on their embeddings. Similar ideas of function approximation have also been widely adopted in MAB literature to share information among large action spaces, with functions ranging from linear ones (Abbasi-Yadkori et al.,, 2011; Yang and Tan,, 2022) to neural networks (Zhu et al.,, 2021; Zhou et al.,, 2020). In the setting considered in this work, we adopt a recently developed BAI-FB scheme as described in the following as TRIPLE-GSE, with details provided in Alg. 4 of Appendix C.
GSE. The general phased elimination flow of SH described in 3.3 is inherited. The major difference is that SH uses sample means to perform eliminations. GSE (Azizi et al.,, 2023), on the other hand, leverages collected samples from previous phases to train a reward function $g_{\bm{\theta}}(\cdot):{\mathbb{R}}^{d}\to{\mathbb{R}}$ that maps prompt embeddings to the predicted performance, which is further used to eliminate prompts.
@@ 5 Experiments
In this section, extensive experimental results are reported to evaluate the efficiency of TRIPLE across diverse prompting tasks from two standard datasets: Instruction-Induction Honovich et al., (2022) and BigBench Suzgun et al., (2022). The results reported in this section are mainly collected from GPT-3.5, Llama2, Gemma, and Mistral (see the specific model numbers listed in Appendix E.1). Full experimental details can be found in Appendix E. The complete results of $47$ tasks are reported in Appendix F, while here we particularly focus on $12$ representative tasks, which are not too hard (i.e., all generated prompts achieve near-zero performances) or too easy (i.e., all generated prompts achieve near-one performances).
@@ 5.1 Evaluating TRIPLE with Fixed Prompt Pools
As TRIPLE main focuses on the prompt selection component, we perform initial evaluations in an isolated fashion of selecting from fixed pools of candidate prompts. For this experiment, candidate pools of prompts are generated following the well-established APE design (Zhou et al.,, 2022) with a high LLM temperature to ensure randomness. Then, under a limited budget, the performances of TRIPLE algorithms are compared with the following four baselines, where the latter two (i.e., BO-EI and NeuralUCB) leverage prompt embeddings:
@@
  • Uniform. Many previous designs choose to evaluate the entire candidate pool on all development data (Guo et al.,, 2023; Prasad et al.,, 2022) which corresponds to uniformly dividing the total budget to test all prompts.
@@
  • UCB. The upper confidence bound (UCB) method is a famous design for regret minimization in MAB. We evaluate UCB using its vanilla version from Auer et al., (2002), which is reported to have good performance in Pryzant et al., (2023).
  • BO-EI. Bayesian optimization (BO) with expected improvement (EI) acquisition function is adopted in Chen et al., (2023), which assumes a Gaussian process prior specified by prompt embeddings to perform posterior updates and makes selection to maximize EI.
  • NeuralUCB. Lin et al., (2023) uses NeuralUCB Zhou et al., (2020) to perform prompt selection, which extends UCB by training a reward function to predict prompt performances based on embeddings.
Table 2: Averaged performance ranks of baselines and TRIPLE on the selected tasks using GPT-3.5, which are computed separately for methods using embeddings or not, where the highest ranked methods are marked bold.
Setup Without embeddings With embeddings
$\bm{|{\mathcal{P}}|}$, $\bm{N}$, LLM Uniform UCB SH CR BO-EI NeuralUCB CLST GSE
$30$, $150$, GPT 3.5 3.28 $\pm$ 0.99 2.50 $\pm$ 1.09 2.04 $\pm$ 1.01 2.29 $\pm$ 1.09 3.67 $\pm$ 0.49 3.00 $\pm$ 1.04 1.68 $\pm$ 0.67 1.60 $\pm$ 0.56
$30$, $150$, Llama2 3.08 $\pm$ 0.79 2.66 $\pm$ 1.07 2.00 $\pm$ 1.27 2.25 $\pm$ 1.13 3.00 $\pm$ 0.95 3.25 $\pm$ 0.75 1.58 $\pm$ 0.67 2.16 $\pm$ 1.26
$30$, $150$, Mistral 3.00 $\pm$ 0.95 2.50 $\pm$ 0.99 2.41 $\pm$ 1.31 2.08 $\pm$ 1.16 3.00 $\pm$ 1.04 2.58 $\pm$ 1.08 2.00 $\pm$ 0.85 2.41 $\pm$ 1.37
$30$, $150$, Gemma 3.21 $\pm$ 1.03 2.46 $\pm$ 1.12 2.04 $\pm$ 1.01 2.29 $\pm$ 1.09 2.91 $\pm$ 0.96 3.16 $\pm$ 1.03 2.04 $\pm$ 1.05 1.87 $\pm$ 0.96
$150$, $100$, GPT 3.5 N/A 2.58 $\pm$ 0.99 3.75 $\pm$ 0.45 2.08 $\pm$ 0.90 1.58 $\pm$ 0.79
$150$, $100$, Llama2 N/A 2.91 $\pm$ 0.95 3.50 $\pm$ 0.65 1.41 $\pm$ 0.49 2.17 $\pm$ 0.98
$150$, $100$, Gemma N/A 2.75 $\pm$ 1.13 3.16 $\pm$ 0.93 2.33 $\pm$ 1.23 1.75 $\pm$ 0.75
@@
Performance with fewer prompts than budget. We first test candidate pools with $30$ prompts per task. Results reported in Fig. LABEL:fig:p_30_N_150_useful reflect the selection performance with an overall budget of $150$. It can be observed that TRIPLE-SH and TRIPLE-CR achieve better performance than Uniform ($15\%$ and $12\%$ improvements on average for GPT-3.5; $15\%$ and $16\%$ for Llama2) and UCB ($5\%$ and $3\%$ improvements on average for GPT-3.5; $6\%$ and $7\%$ for Llama2). Moreover, for methods using prompt embeddings, the enhanced TRIPLE-CLST and TRIPLE-GSE also demonstrate remarkable improvements over BO-EI ($11\%$ and $10\%$ on average for GPT-3.5; $56\%$ and $52\%$ for Llama2) and NeuralUCB ($17\%$ and $17\%$ improvements on average for GPT-3.5; $26\%$ and $27\%$ for Llama2). These results empirically evidence the superiority of TRIPLE with or without prompt embeddings.
Performance with more prompts than budget. In the above test, the budget is larger than the number of candidate prompts. We further perform experiments in a more difficult setting, i.e., there are more prompts than the budget. In particular, candidate pools with $150$ prompts per task are generated, and the overall budget is set as $100$. In this scenario, only the methods that can leverage embeddings (i.e., BO-EI, NeuralUCB, TRIPLE-CLST, TRIPLE-GSE) can be used, as otherwise the total budget is not sufficient to provide even one evaluation to initiate the performance estimation of each candidate prompt. Results are reported in Fig. LABEL:fig:p_150_N_100_useful. In particular, it can be observed that TRIPLE-CLST and TRIPLE-GSE significantly improve over BO-EI ($21\%$ and $28\%$ on average for GPT-3.5; $31\%$ and $42\%$ for Llama2) and NeuralUCB ($38\%$ and $45\%$ on average for GPT-3.5; $26\%$ and $16\%$ for Llama2).
A summary of the averaged performance ranks of the baselines and TRIPLE is listed in Table 2, which contains results on four LLMs (i.e., GPT-3.5, Llama2, Mistral, Gemma). It can be observed that in varying setups and with different LLMs, the proposed TRIPLE methods consistently obtain better performances than the previous baselines, remarking its efficiency and broad applicability.
Figure 3: Relative gain of over Uniform under different budgets, collected with GPT-3.5.
@@
Impact of the total budget. For a more comprehensive understanding, using candidate pools with $30$ prompts, we further examine the impact of budgets, starting with $5$ evaluations per prompt on average (i.e., $150$ overall as adopted in Fig. LABEL:fig:p_30_N_150_useful), and then gradually increasing to $30$ (i.e., $900$ overall, which is the same as the experiments in Zhou et al., (2022)). From the results shown in Fig. 3, we see that the improvements of TRIPLE over baselines are more pronounced with lower budgets. In particular, with a budget of $10$ evaluations per prompt on average, TRIPLE-CR, TRIPLE-CLST and TRIPLE-GSE maintain notable $9.7\%$, $13.5\%$ and $17.4\%$ improvement over Uniform, respectively; when the budget escalates to $20$ evaluations per prompt on average, TRIPLE-CLST and TRIPLE-GSE still achieve an approximate $8\%$ improvement. Once the budget reaches $30$ evaluations per prompt on average, all methods provide approximately the same performance as they can all identify the optimal prompts under this generous budget.
@@
Table 3: Performances of integrating TRIPLE in the end-to-end pipelines using GPT-3.5. The baseline methods reported in the original implementations are labeled as (b). For each task, the best score across two pipelines is marked as red, and the best score in the remaining pipeline is highlighted as yellow. TRIPLE-CR are selected over TRIPLE-SH due to its better performance observed in the previous experiments. TRILE-CLST is ignored in the tests with APO, as it is ineffective to cluster only $10$ prompts.